Find the distance between the two points.

M(1,-3),N(3,5)

thank you

Answer by Alan3354(69443)

Find the distance between the two points.
M(1,-3),N(3,5)
----------
use Pythagoras.
d^2 = dffy^2 = diffx^2 = (3-1)^2 + (5+3)^2
d^2 = 4 + 64 = 68
d = sqrt(68) = 2sqrt(17)